# Multi-Hypothesis Prediction for Portfolio Optimization: A Structured Ensemble Learning Approach to Risk Diversification

### v5

@@ 4.4 Comparative Performance Analysis
This section presents a comparative analysis between the proposed method, implemented with the s-RBFN as the PSEM model, and several well-established robust and data-driven ”plug-in” portfolio optimization methods, including Inverse Volatility (IV) (ANG et al., 2006), CVaR Risk Parity (CVaR RP) (Kapsos et al., 2018), Maximum Diversification (MD) (Choueifaty and Coignard, 2008), Hierarchical Risk Parity (HRP) (Lopez de Prado, 2016) and Hierarchical Equal Risk Contribution (HERC) (Raffinot, 2018). The comparison is conducted over a two-year investment horizon with monthly portfolio reallocation. The performance metric used is the monthly standard Sharpe Ratio, calculated using a time series of 1-month cumulative returns. This is computed as the average return over the 24-month period divided by its standard deviation and then annualized. It is important to note that the Sharpe ratios reported in this section are generally lower because they are calculated using average returns, whereas in previous sections a modified version of the Sharpe ratio was used, based on cumulative returns over 1-month and 10-month periods.
Table 4 reports the average Sharpe Ratios across various threshold levels $T$ and values of the asset selection diversity parameter $\gamma$. It can be shown how the presented method to select assets based predictions diversity (diversify-quality trade-off of predictions as in Section 4.3.2) makes all the methods improve generalization performance due to increased out-of-sample diversification measured by the Sharpe ratios. Also, the s-RBFN method is the best as can be seen in the average value across all the experiments in the same table at the bottom. For selections of assets with more than -0.5% cumulative monthly return as in Table 4(a), diversity is not of clear value but for selections above 0.0% and 0.3% it is a clear wining selection choice for all methods.
@@
Figure 9 displays the distribution of Sharpe ratios, ordered from highest to lowest, across all experiments conducted in this section (108 in total). These experiments include various configurations of the diversity parameters $\varepsilon$ and $\gamma$, threshold $T$, and the number of portfolio constituents $M$. The horizontal axis represents individual experiments, sorted by decreasing Sharpe ratio. The plot clearly demonstrates that the s-RBFN consistently outperforms all other models—often by a significant margin—highlighting the robustness and accuracy of the proposed method.
@@ 5 Conclusion and Future Work
This work explored the relationship between the diversity of individual predictors in structured ensembles and portfolio diversification, offering a novel perspective within a multiple-hypothesis predict-then-optimize framework. One key contribution was the use of the ensemble combiner rule as the portfolio target during optimization—for example, the equal-weighted portfolio with the MSE loss. To the best of the authors’ knowledge, this represents the first plug-in method that enables parametric control of out-of-sample diversification prior to decision-making. Experimental results confirmed that out-of-sample diversification and generalization performance depend not only on the optimization procedure but also on the asset selection stage. Notably, when RMSE ranges were similar, asset selections with greater diversity in return predictions consistently outperformed those with less diversity, highlighting the importance of the diversity-quality trade-off. This forms a third contribution, and both sources of diversity were shown to extend the boundaries of portfolio diversification beyond previously suggested limits (Dalio, 2018).
@@ Appendix B Ensemble Regularization in the Gaussian Basis Functions Experiments (Multi-step ahead case)
To perform a closer analysis of the regularization parameter, it can be seen how for $\lambda_{s}=10$ the generalization error is more stable for all model hyperparameters, and in the case of no regularization parameter, there are some cases in which the results are quite unstable (like in Figures 12(b) and 12(c)) were the test RMSE are displayed. It can be concluded that the regularization parameter reduces the uncertainty of the s-RBFN hyperparameters in the case of the Gaussian basis function. This instability in hyperparameter selection was not observed with the s-RBFN using radial basis functions.

### v6

@@ 4.4 Comparative Performance Analysis
This section presents a comparative analysis between the proposed method, implemented with the s-RBFN as the PSEM model, and several well-established, robust and data-driven ”plug-in” portfolio optimization methods. These consist of Inverse Volatility (IV) (ANG et al., 2006), CVaR Risk Parity (CVaR RP) (Kapsos et al., 2018), Maximum Diversification (MD) (Choueifaty and Coignard, 2008), Hierarchical Risk Parity (HRP) (Lopez de Prado, 2016) and Hierarchical Equal Risk Contribution (HERC) (Raffinot, 2018). The comparison is conducted over a two-year investment horizon with monthly portfolio reallocation. The performance metric used is the monthly standard Sharpe Ratio, calculated using a time series of 1-month cumulative returns. This is computed as the average return over the 24-month period divided by its standard deviation and then annualized. It is important to note that the Sharpe ratios reported in this section are generally lower because they are calculated using average returns, whereas in previous sections, a modified version of the Sharpe ratio was used, based on cumulative returns over 1-month and 10-month periods.
Table 4 reports the average Sharpe Ratios across various threshold levels $T$ and values of the asset selection diversity parameter $\gamma$. It can be shown how the presented method to select assets based on prediction diversity (diversify-quality trade-off of predictions as in Section 4.3.2) makes all the methods improve generalization performance, due to increased out-of-sample diversification measured by the Sharpe ratios. Also, the s-RBFN is the top-performing method, as can be seen in the average value across all the experiments at the bottom of the same table. For selections of assets with less than -0.5% cumulative monthly return, as in Table 4(a), diversity is not of clear value, but for selections above 0.0% and 0.3% it is a clear winning selection choice for all methods.
@@
Figure 9 displays the distribution of Sharpe ratios, ordered from highest to lowest, across all experiments conducted in this section (108 in total). These experiments include various configurations of the diversity parameters $\varepsilon$ and $\gamma$, threshold $T$, and the number of portfolio constituents $M$. The horizontal axis represents individual experiments, sorted by decreasing Sharpe ratio. Figure 9 clearly demonstrates that the s-RBFN consistently outperforms all other models—often by a significant margin—highlighting the robustness and accuracy of the proposed method.
@@
In D.2, the proposed method is compared against alternative portfolio optimization approaches across distinct market regimes—specifically, the bear market during the Global Financial Crisis (2007–2009) and the bull market of the subsequent U.S. equity rally (2009–2016). Additionally, a diversified global bond data set comprising 1,336 instruments from 2014 to 2018, characterized in Figure 15, is used to evaluate performance across varying portfolio sizes.
As shown in Tables 8 and 9, the proposed s-RBFN model consistently achieves the highest average performance across all portfolio sizes under both bull and bear market regimes. Table 10 presents the results for the fixed income data set, where the s-RBFN attains superior Sharpe, Sortino, and Omega ratios, along with the second-lowest maximum drawdown among all evaluated methods. Notably, the s-RBFN model with 10 portfolio constituents delivers the best performance across all metrics for both the U.S. equity portfolios during the Credit Crunch (Table 9) and the fixed income data set (Table 10).
Overall, the proposed solution demonstrates significantly stronger performance for small-sized portfolios compared to other methods, which can be attributed to the enhanced diversification induced by the diversity parameters.
@@ 5 Conclusion and Future Work
This work explored the relationship between the diversity of individual predictors in structured ensembles and portfolio diversification, offering a novel perspective within a multiple-hypothesis predict-then-optimize framework. One key contribution was the use of the ensemble combiner rule as the portfolio target during optimization—for example, the equal-weighted portfolio with the MSE loss. To the best of the authors’ knowledge, this represents the first learning-based plug-in method that enables parametric control of out-of-sample diversification prior to decision-making. Experimental results confirmed that out-of-sample diversification and generalization performance depend not only on the optimization procedure but also on the asset selection stage. Notably, when RMSE ranges were similar, asset selections with greater diversity in return predictions consistently outperformed those with less diversity, highlighting the importance of the diversity-quality trade-off. This forms a third contribution, and both sources of diversity were shown to extend the boundaries of portfolio diversification beyond previously suggested limits (Dalio, 2018).
@@
Acknowledgements
The authors gratefully acknowledge Miraltabank Asset Management, and in particular CIO Ignacio Fuertes Aguirre, for providing the fixed income data used in the experiments.
@@ Appendix B Ensemble Regularization in the Gaussian Basis Functions Experiments (Multi-step ahead case)
To perform a closer analysis of the regularization parameter, it can be seen how for $\lambda_{s}=10$ the generalization error is more stable for all model hyperparameters, and in the case of no regularization parameter, there are some cases in which the results are quite unstable (like in Figures 12(b) and 12(c)) where the test RMSE are displayed. It can be concluded that the regularization parameter reduces the uncertainty of the s-RBFN hyperparameters in the case of the Gaussian basis function. This instability in hyperparameter selection was not observed with the s-RBFN using radial basis functions.
@@ Appendix C Including Diversity in the Asset or Hypothesis Selection Stage: One-step Decision-making
Appendix D Sensitivity Analysis of Diversity Parameters across different Regimes, Asset Classes and Performance Metrics
D.1 Performance Analysis of the s-RBFN Model Across Different Regimes and Asset Classes as a Function of the Diversity Parameter $\varepsilon$ and the Number of Portfolio Constituents $n$
Table 5: Performance metrics across varying values of $\varepsilon$ and number of assets $n$ during the U.S. equities market rally spanning 2009 to 2016.
Metric $n$ 0 0.1 0.35 0.5 0.75 0.95
Sharpe 5 0.645 0.629 0.399 1.035 0.685 0.691
10 1.093 1.208 0.960 0.687 1.023 1.220
20 1.369 1.135 1.135 1.452 1.135 0.852
100 1.020 1.182 1.061 1.231 1.229 1.186
Avg. 1.032 1.039 0.889 1.101 1.018 0.987
Max Drawdown 5 -0.307 -0.286 -0.219 -0.328 -0.336 -0.285
10 -0.352 -0.252 -0.258 -0.348 -0.322 -0.215
20 -0.261 -0.235 -0.220 -0.259 -0.248 -0.289
100 -0.260 -0.259 -0.261 -0.231 -0.258 -0.252
Avg. -0.295 -0.258 -0.240 -0.292 -0.291 -0.260
Sortino 5 0.923 0.871 0.536 1.417 0.893 0.983
10 1.403 1.626 1.336 0.934 1.397 1.763
20 1.897 1.532 1.541 1.900 1.453 1.125
100 1.328 1.541 1.385 1.520 1.611 1.557
Avg. 1.388 1.393 1.200 1.443 1.339 1.357
Omega 5 1.123 1.118 1.073 1.205 1.130 1.132
10 1.221 1.242 1.185 1.129 1.198 1.237
20 1.270 1.224 1.224 1.296 1.211 1.165
100 1.201 1.238 1.211 1.231 1.246 1.237
Avg. 1.204 1.206 1.173 1.215 1.196 1.193
Table 6: Performance metrics across varying values of the diversity parameter $\varepsilon$ and number of assets $n$, averaged over $M$, during the U.S. Credit Crunch Crisis (2007–2009).
Metric $n$ 0 0.1 0.35 0.5 0.75 0.95
Sharpe 5 -0.454 0.007 0.380 -0.027 -0.744 0.585
10 -0.457 -0.027 -0.225 -0.203 -0.484 -0.112
20 0.129 -0.041 -0.034 0.215 -0.284 -0.226
100 -0.186 -0.170 -0.209 -0.124 -0.264 -0.298
Avg. -0.242 -0.058 -0.022 -0.035 -0.429 -0.013
Max Drawdown 5 -0.646 -0.690 -0.605 -0.645 -0.790 -0.488
10 -0.713 -0.564 -0.604 -0.607 -0.640 -0.534
20 -0.498 -0.570 -0.551 -0.382 -0.578 -0.575
100 -0.532 -0.540 -0.577 -0.545 -0.581 -0.606
Avg. -0.597 -0.591 -0.584 -0.545 -0.647 -0.551
Sortino 5 -0.623 0.008 0.542 -0.036 -1.077 0.846
10 -0.625 -0.038 -0.319 -0.270 -0.749 -0.167
20 0.191 -0.053 -0.047 0.309 -0.401 -0.313
100 -0.267 -0.238 -0.291 -0.173 -0.372 -0.412
Avg. -0.331 -0.080 -0.029 -0.043 -0.900 -0.011
Omega 5 0.921 1.001 1.073 0.995 0.878 1.111
10 0.912 0.995 0.960 0.964 0.916 0.981
20 1.024 0.992 0.994 1.039 0.949 0.960
100 0.967 0.970 0.963 0.978 0.954 0.947
Avg. 0.956 0.990 0.998 0.994 0.924 1.000
(a) Rating
(b) Tenor (curve)
(c) Seniority
(d) Country (ISO)
(e) Sector
Figure 15: Distribution of the 1,336 bonds in the data set across key categories: rating, curve (tenor), seniority, country (ISO), and sector.
Table 7: Performance metrics across varying values of $\varepsilon$ and number of assets $n$ on a fixed income data set covering the global bond market (public, private, institutional, corporate, municipal sectors) from 2014 to 2018.
Metric $n$ 0 0.1 0.35 0.5 0.75 0.95
Sharpe 5 -0.25 -0.28 0.01 -0.68 -0.31 0.03
10 0.48 -0.57 -0.20 0.51 0.15 -0.44
20 -0.31 0.51 -0.54 -0.55 -0.33 -0.20
100 0.41 0.31 0.52 -0.48 0.39 0.35
Avg. 0.08 -0.01 -0.05 -0.30 -0.02 -0.06
Max Drawdown 5 -0.10 -0.21 -0.07 -0.19 -0.17 -0.09
10 -0.14 -0.10 -0.06 -0.15 -0.07 -0.10
20 -0.06 -0.12 -0.10 -0.08 -0.09 -0.05
100 -0.08 -0.06 -0.06 -0.07 -0.05 -0.06
Avg. -0.10 -0.12 -0.07 -0.12 -0.10 -0.08
Sortino 5 -0.29 -0.29 0.01 -0.75 -0.32 0.04
10 4.52 -0.67 -0.24 9.34 0.18 -0.49
20 -0.36 12.24 -0.61 -0.64 -0.43 -0.24
100 2.97 1.36 16.16 -0.62 2.17 1.17
Avg. 1.21 3.16 3.33 1.83 0.40 0.12
Omega 5 0.93 0.91 1.00 0.84 0.91 1.01
10 2.84 0.88 0.96 5.01 1.03 0.90
20 0.94 4.76 0.89 0.90 0.91 0.96
100 1.44 1.20 3.79 0.91 1.31 1.18
Avg. 1.54 1.94 1.66 1.91 1.29 1.26
D.2 Comparative Analysis of Different Portfolio Optimization Methods
Table 8: Performance metrics by strategy and number of assets during the U.S. Equity Rally regime.
n s-RBFN 1_N CVaR RP HERC HRP IV MD
Sharpe Ratio
5 1.11 0.51 0.58 0.08 0.48 0.58 0.49
10 1.03 1.11 1.10 1.20 1.17 1.10 1.16
20 1.16 1.12 1.20 1.00 1.18 1.20 1.11
100 1.15 1.13 1.19 0.97 1.23 1.19 0.83
Avg 1.11 0.97 1.02 0.81 1.02 1.02 0.90
Maximum Drawdown
5 -0.30 -0.31 -0.28 -0.41 -0.29 -0.28 -0.26
10 -0.31 -0.34 -0.27 -0.27 -0.25 -0.27 -0.27
20 -0.26 -0.22 -0.20 -0.23 -0.20 -0.20 -0.19
100 -0.25 -0.26 -0.21 -0.23 -0.21 -0.21 -0.32
Avg -0.28 -0.28 -0.24 -0.29 -0.24 -0.24 -0.26
Sortino Ratio
5 1.25 0.71 0.80 0.11 0.65 0.80 0.68
10 1.46 1.39 1.41 1.67 1.52 1.41 1.51
20 1.64 1.50 1.61 1.26 1.57 1.61 1.52
100 1.55 1.47 1.53 1.26 1.58 1.53 1.07
Avg 1.48 1.27 1.34 1.08 1.33 1.34 1.20
Omega Ratio
5 1.00 0.94 0.94 0.98 0.94 0.94 0.94
10 1.17 0.91 0.90 0.92 0.89 0.90 0.87
20 1.02 0.96 0.95 0.93 0.94 0.95 0.93
100 0.99 0.97 0.95 0.95 0.95 0.95 0.95
Avg 1.04 0.95 0.93 0.95 0.93 0.93 0.92
Table 9: Performance metrics by strategy and number of assets during the U.S. Credit Crunch (2007–2009).
Assets s-RBFN 1/N CVaR RP HERC HRP IV MD
Sharpe Ratio
5 -0.64 -0.33 -0.37 -0.11 -0.34 -0.37 -0.35
10 0.82 -0.53 -0.59 -0.42 -0.65 -0.59 -0.72
20 -0.15 -0.22 -0.28 -0.38 -0.34 -0.28 -0.38
100 -0.07 -0.18 -0.28 -0.23 -0.31 -0.28 -0.28
Avg. -0.01 -0.32 -0.38 -0.29 -0.41 -0.38 -0.43
Maximum Drawdown
5 -0.80 -0.60 -0.57 -0.54 -0.58 -0.57 -0.56
10 -0.60 -0.68 -0.61 -0.69 -0.62 -0.61 -0.62
20 -0.64 -0.55 -0.52 -0.59 -0.54 -0.52 -0.53
100 -0.56 -0.53 -0.51 -0.50 -0.51 -0.51 -0.48
Avg. -0.65 -0.59 -0.55 -0.58 -0.56 -0.55 -0.55
Sortino Ratio
5 -0.12 -0.49 -0.55 -0.17 -0.50 -0.55 -0.52
10 0.07 -0.72 -0.80 -0.52 -0.86 -0.80 -0.91
20 -0.21 -0.31 -0.39 -0.50 -0.48 -0.39 -0.50
100 -0.11 -0.25 -0.40 -0.31 -0.44 -0.40 -0.37
Avg. -0.09 -0.44 -0.54 -0.38 -0.57 -0.54 -0.58
Omega Ratio
5 0.89 0.94 0.94 0.98 0.94 0.94 0.94
10 1.01 0.91 0.90 0.92 0.89 0.90 0.87
20 0.97 0.96 0.95 0.93 0.94 0.95 0.93
100 0.99 0.97 0.95 0.95 0.95 0.95 0.95
Avg. 0.97 0.95 0.94 0.95 0.93 0.94 0.92
Table 10: Performance metrics by strategy and number of assets on a data set of 1,336 global bonds, including public, corporate, institutional, municipal, and other sectors (2014–2018).
Assets s-RBFN 1/N IV MD HRP HERC
Sharpe Ratio
5 -0.25 -0.48 -0.80 -0.65 -0.73 -0.78
10 0.48 0.47 0.29 -0.14 0.32 0.19
20 -0.31 -0.37 -0.42 0.08 -0.16 -0.37
100 0.41 0.39 0.32 0.06 0.36 -0.02
Avg. 0.08 0.00 -0.15 -0.16 -0.05 -0.25
Maximum Drawdown
5 -0.10 -0.11 -0.12 -0.10 -0.11 -0.12
10 -0.14 -0.12 -0.12 -0.13 -0.13 -0.14
20 -0.06 -0.06 -0.09 -0.17 -0.11 -0.09
100 -0.08 -0.08 -0.15 -0.07 -0.12 -0.18
Avg. -0.10 -0.09 -0.12 -0.12 -0.12 -0.13
Sortino Ratio
5 -0.29 -0.54 -0.73 -0.59 -0.67 -0.73
10 4.52 4.42 1.91 -0.25 2.19 0.80
20 -0.36 -0.48 -0.58 0.11 -0.23 -0.50
100 2.97 2.85 4.09 0.09 3.87 -0.13
Avg. 1.21 1.06 1.17 -0.16 1.29 0.11
Omega Ratio
5 0.93 0.90 0.80 0.82 0.81 0.81
10 2.84 2.62 1.53 0.91 1.69 1.20
20 0.94 0.91 0.79 1.11 0.88 0.84
100 1.44 1.40 1.65 1.07 1.79 0.98
Avg. 1.53 1.46 1.19 0.98 1.29 0.96